OpenStudy (anonymous):

tennis ball has a diameter of approximately 6.5 centimeters, what is its surface area?

OpenStudy (anonymous):

shere SA= 4(pi)r^2 so this is simple. we get SA=4(pi)(6.5)^2 which comes out to 169(pi)

OpenStudy (anonymous):

make sure to include the squared units! like feet squared or meters squared!